OpenStudy (anonymous):

Help Please. I'm going out of my mind. I thought this seemed pretty straight forward, but I have tried it literally 17 times and I cannot for the life of me figure it out. help me please. here is the problem: f(x) = x^2+2x+3 and g(x) = 4x^3-5x^2+4x+1 For brevity let's write deg (p) for the degree of a polynomial p. So in the above example, deg (f) = 2 and deg (g) = 3. You can answer the following questions without actually computing the indicated functions. deg (f+g) . deg (f-g) . deg (fg) . deg (f o g) . deg (g o f).

OpenStudy (perl):

deg (f+g) = max deg (f,g) deg(f-g) = max deg (f,g) deg (f*g) = deg(f) + deg (g)

OpenStudy (anonymous):

deg (f+g) is 3 deg (f-g) . is 3 deg (fg) . is 6

OpenStudy (anonymous):

oops deg(fg) is 5

OpenStudy (anonymous):

deg of (fog) is 6 deg (g o f) is 6

OpenStudy (perl):

deg (f +/- g) = max {deg (f), deg(g)}

OpenStudy (anonymous):

how are you getting that without having to compute?

OpenStudy (perl):

you can ignore the other terms, the dominant terms are the highest degree

OpenStudy (anonymous):

ohh! I think i was over thinking it and over computing. thank you so much!

OpenStudy (anonymous):

follow the rules as defined by perl

OpenStudy (anonymous):

@perl i wrote for f o g or gof but next i hv corrected it

OpenStudy (perl):

deg (f+g) = max {deg (f), deg(g)} deg(f-g) = max {deg (f) deg,(g)} deg (f*g) = deg(f) + deg (g) deg ( f o g ) = deg (f) * deg(g) deg (gof) = deg (f) * deg (g) we should prove this, but shouldnt be hard basically i got the formulas from just manipulating the dominant terms

OpenStudy (perl):

@matricked that was correct, the deg(f*g) was 5, but you corrected

OpenStudy (anonymous):

for f+g the dominate degrees are 2 and 3, wouldn't that make the answer 5 and not 3?

OpenStudy (perl):

i have to make some conditions on my rules Theorem 1.1 Let R be a ring and f, g ∈ R[x] {0}. The following statements are true.

OpenStudy (perl):

(i) fg = 0 or deg (fg) ≤ deg f + deg g. (ii) f + g = 0 or deg (f + g) ≤ max{deg f, deg g}. (iii) If deg f = deg g, then deg(f + g) = max{deg f, deg g}. (iv) fg = 0 or ord (fg) ≥ ord f + ord g. (v) f + g = 0 or ord (f + g) ≥ min{ord f, ord g}. (vi) If ord f = ord g, then ord(f + g) = min{ord f, ord g}.

OpenStudy (anonymous):

wow. I'm trying to study this for my exam in about 15 days but i fell confidant that I will bomb any problem like this

OpenStudy (perl):

I would just manipulate the dominant terms ,when you multiply the two expression, the other terms are all smaller

OpenStudy (perl):

( x^2+2x+3) (4x^3-5x^2+4x+1) , for example if you multiply them you get ------------------------ 4x^5 + smaller power terms

OpenStudy (perl):

so the degree of the product of the polynomials is going to be 5

OpenStudy (anonymous):

oh, I think I see what you are saying.

OpenStudy (perl):

assuming you put the polynomials in descending order (standard form)

OpenStudy (perl):

the only tricky one is the composition

OpenStudy (perl):

and maybe division, but you werent asked

OpenStudy (anonymous):

Thank you. I'll probably be studying this and trying to understand it, still feels foreign.

OpenStudy (perl):

:)

OpenStudy (perl):

try the composition , using the idea that you only need the dominant terms

OpenStudy (perl):

fog = f(g(x)) = f ( 4x^3 + smaller terms ) = (4x^3 + smaller terms ) ^2 = (4x^3) ^2 + smaller terms = 16x^6 + smaller terms, and the degree is 6

OpenStudy (anonymous):

ohmigawd....i think i really understood what you just wrote! Thank you jezus! seriously, thank you for your help!

OpenStudy (perl):

:)

OpenStudy (perl):

finding general rules can be tricky, since there are different cases to consider. so you can ignore the rules i wrote above, i have to double check them

OpenStudy (anonymous):

okay

OpenStudy (perl):

and just use the trick of looking at the dominant terms plus smaller terms

OpenStudy (anonymous):

I will thank you.